Following is information on two alternative investments being considered by Jolee Company. The company requires a 8% return from its investments. (PV of $1, FV of $1, PVA of $1, and FVA of $1) (Use appropriate factor(s) from the tables provided.)

Project A Project B
Initial investment $ (185,325 ) $ (160,960 )
Expected net cash flows in:
Year 1 52,000 38,000
Year 2 53,000 51,000
Year 3 74,295 63,000
Year 4 92,400 70,000
Year 5 58,000 21,000


a. For each alternative project compute the net present value.
b. For each alternative project compute the profitability index. If the company can only select one project, which should it choose?

Complete this question by entering your answers in the tabs below.

  • Required A
  • Required B

For each alternative project compute the net present value.

Project A
Initial Investment $185,325
Chart Values are Based on:
i = %
Year Cash Inflow x PV Factor = Present Value
1 =
2 =
3 =
4 =
5 =
Project B
Initial Investment $160,960
Year Cash Inflow x PV Factor = Present Value
1 =
2 =
3 =
4 =
5 =
  • Required B

For each alternative project compute the profitability index. If the company can only select one project, which should it choose?

Profitability Index
Choose Numerator: / Choose Denominator: = Profitability Index
/ = Profitability index
Project A 0
Project B 0
If the company can only select one project, which should it choose?
Answer #1

Solution a:

Project A
Initial investment $185,325.00
Chart values are based on:
i= 8%
Year Cash inflows * PV Factor = Present Value
1 $52,000.00 * 0.92593 = $48,148
2 $53,000.00 * 0.85734 = $45,439
3 $74,295.00 * 0.79383 = $58,978
4 $92,400.00 * 0.73503 = $67,917
5 $58,000.00 * 0.68058 = $39,474
Total $259,955
Present value of cash inflows $259,955
Initial investment $185,325
Net Present Value $74,630
Project B
Initial investment $160,960.00
Chart values are based on:
i= 8%
Year Cash inflows * PV Factor = Present Value
1 $38,000.00 * 0.92593 = $35,185
2 $51,000.00 * 0.85734 = $43,724
3 $63,000.00 * 0.79383 = $50,011
4 $70,000.00 * 0.73503 = $51,452
5 $21,000.00 * 0.68058 = $14,292
Total $194,665
Present value of cash inflows $194,665
Initial investment $160,960
Net Present Value $33,705

Solution b:

Profitability Index
Choose Numerator / Choose Denominator = Profitability Index
NPV / Initial investment = Profitability Index
Project A $74,630 / $185,325 = 0.40
Project B $33,705 / $160,960 = 0.21
If the company can select one project, which should it choose? Project A